OpenStudy (anonymous):

Given f(x) = the quantity of 4x plus 1, divided by 3, solve for f-1(3).

OpenStudy (anonymous):

\[f(x)=\frac{4x+1}{3}\]?

OpenStudy (anonymous):

and you want \(f^{-1}(3)\) right?

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

interpret this question as an algebra question: solve for \(x\) \[\frac{4x+1}{3}=3\]

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

x=2??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

this means \(f(2)=3\) so \(f^{-1}(3)=2\)

OpenStudy (anonymous):

yw